OpenStudy (anonymous):

HELP IN 6th GRADE NO CALCULATORS what is 45/5= A. 9 R 5 B. 9 R 4 C. 10 D. None of above

OpenStudy (bakonloverk):

You dont need a calculator

OpenStudy (bakonloverk):

Do it the old fashioned way, paper and pencil.

OpenStudy (anonymous):

here is the A. 9 R 5 B. 9 R 4 C. 10 D. None of above

OpenStudy (anonymous):

i think it is B am i right

OpenStudy (bakonloverk):

wait 5/49 or 49/5?

OpenStudy (solomonzelman):

it seems to me as though as you don't have enough background knowledge for this question.

OpenStudy (♪chibiterasu):

I remember when I used paper and pencil as a children, I was so terrible that I asked my parents to change my name so Loser McGee. No joke, I as a stupid kid. But yeah, 5/49 is a really long decimal, so 49/5 would be a little more logical . . .

OpenStudy (anonymous):

it is 45/5

OpenStudy (♪chibiterasu):

Oh, then that's very easy.

OpenStudy (bakonloverk):

Yes, very easy.

OpenStudy (anonymous):

i think it is B am i Right?

OpenStudy (solomonzelman):

The other way \[\frac{45}{5}\] Divide top and bottom by 5. \[\frac{45\color{red} {\div 5}}{5\color{red} {\div 5}}=?\]

OpenStudy (anonymous):

45/5=9 5/5=1

OpenStudy (♪chibiterasu):

Let's to it the old fashioned way in kinder: There is [insert answer here] in each bucket of magical doodoo. Now start counting with your fingers.

OpenStudy (♪chibiterasu):

Lol the first bucket has one less than it should dang it

OpenStudy (anonymous):

45/5=9 5/5=1

OpenStudy (♪chibiterasu):

Yeah 9 in each doodoo bucket, gratz :P

OpenStudy (solomonzelman):

\[\frac{45\color{red} {\div 5}}{5\color{red} {\div 5}}=\frac{?}{?}=?\]FILL IN BLANKS

OpenStudy (♪chibiterasu):

9/1 is the same as just 9. SZ he already did that ;-;

OpenStudy (♪chibiterasu):

But none of the options are a whole 9 so pretty much D.

OpenStudy (bakonloverk):

yeah D

OpenStudy (anonymous):

45/5= A. 9 R 5 B. 9 R 4 C. 10 D. None of above

OpenStudy (anonymous):

so D

OpenStudy (♪chibiterasu):

D, since no options are a whole 9.

OpenStudy (♪chibiterasu):

Yeah :P

OpenStudy (bakonloverk):

Comon Sense says the answer is D

OpenStudy (♪chibiterasu):

This math book has problems ಠ_ಠ

OpenStudy (anonymous):

thanks can you help with one more please

OpenStudy (♪chibiterasu):

Sure ಠ_ಠ

OpenStudy (anonymous):

8/352 A. 44 B. 46 C. 43 D. 344

OpenStudy (anonymous):

D

OpenStudy (anonymous):

well that is it thanks

OpenStudy (♪chibiterasu):

Ummmm that's a really long decimal again.

OpenStudy (bakonloverk):

8/352 is a decimal. 352/8 is 44

OpenStudy (♪chibiterasu):

Lol are you typing these in right? :P

OpenStudy (bakonloverk):

Thats what i was thinking

OpenStudy (anonymous):

i passed my test thanks Guys :D

OpenStudy (♪chibiterasu):

ಠ_ಠ

OpenStudy (bakonloverk):

woah woah woah, this was a test?

OpenStudy (anonymous):

the answer is just D. none of the above. its 6th grade math dont over think it.

OpenStudy (♪chibiterasu):

I read "test" and I think hmm ಠ_ಠ

OpenStudy (bakonloverk):

exactly

OpenStudy (anonymous):

nope this was a pre test got u guys

OpenStudy (♪chibiterasu):

Lol we're criminals now RUN

OpenStudy (bakonloverk):

Shame, if it was a pretest you obviously dont know how to divide

OpenStudy (♪chibiterasu):

Is this FLVS? Because pre-tests are just worth 1 point of your grade so I guess on all of them.

OpenStudy (anonymous):

LOL :D

OpenStudy (bakonloverk):

and i mean that in the nicest way possible

OpenStudy (anonymous):

ya right

OpenStudy (♪chibiterasu):

Well in pre-tests they test what you already know so not as much shame but still some shame ಠ_ಠ

OpenStudy (anonymous):

well bye thanks for help i think

OpenStudy (anonymous):

can i fan u bakonloverk

OpenStudy (bakonloverk):

Sure?

OpenStudy (anonymous):

thanks bye :D

OpenStudy (♪chibiterasu):

Lol I though I already had bakon fanned lel :P Lag tho man ಠ_ಠ

OpenStudy (bakonloverk):

hmm i thought you did too